OpenStudy (anonymous):

i have a math problem.... : 2r+6 over 3+2r-1 over 2 = 7r over 6-(r-4)

OpenStudy (anonymous):

\[\frac{\frac{2r+6}{3+2r-1}}{\frac{2+7r}{6-(r-4)}}\] Is this correct?

OpenStudy (hoblos):

oh whats that??!! i dont think so

OpenStudy (hoblos):

(2r+6)/3 + (2r-1)/2 = 7r/6 - (r-4) it it like this heather49 or what